answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Which term describes the set of all possible outcomes of an experiment?

  1. if you were to take a quarter and flip it, Whats the probability that it lands on head. The answer would be 1/2 because theirs only two sides two have a fifty fifty chance on landing to either side.
  2. Bad example: There is a four sided rubrics-cube, you wanna make it land on the red color, well that my lilt friend would be 1/2.

How many diamonds are in a deck?

There Are 13 Card Of each Suit.. Total There are 52 card and 4 suits.. On Calculating We get 13 Cards For Each Suit..

Why do univariate distributions matter in terms of social science and statistics why should we graph univariate distribution before we numerically summarize them?

Not all univariate data will be normally distributed. Graphing the data will help you determine if you got the kind of distribution you were expecting, and if not, what kinds of tests will be appropriate for what you got. A strange distribution when you had reason to expect, say, a normal distribution would help you uncover possible problems with data collection.

How many heads can you have?

The law of large numbers infers you will have 50% heads.

If dealt one card from a standard 52 card deck what is the probability of being dealt an ace or a 9?

Aces and 9s are disjoint events, so the probability of either is the sum of the probabilities of each.

P(A or 9) = P(A) + P(9) = 1/13 + 1/13 = 2/13

What is the probability that both of them get 'selected is Bruce and Jack apply for a job in the same company. The chance of Bruce getting selected is 25 and that of Jack is 47.?

Assuming you mean probabilities of 0.25 and 0.47, multiply them together to get the combined probabilities.This assumes that the probabilities are independent of one another. This is not always the case in real life; for example, if the company only has one open position, the probability of both being selected would be ZERO.

What is the probability that the card is a six?

In an ordinary deck of 52 cards, the probability that a single, randomly drawn card is a six, is 1/13.

What is a large event called?

Actually there are four names: Graduation, wedding, pregnancy, divorce. Death used to be one, but then it became inconsequential due to the lack of interest by the participant.

What does 'not mutually exclusive of one another' mean?

If two events ARE mutually exclusive, then it means that they can not both happen simultaneously. For example, if we flip a coin, it can only be heads or tails, not both. an example of not mutually exclusive events are strong winds and rain. it could be strong wind, or rain, or both.

What is alike between experimental and theoretical probability?

They are probabilities: that is, estimates of the likelihood of an event happening.

What is he probability that a spinner will land on A if there are two A's on the spinner?

It depends on how many other positions are on the spinner. The question, as asked, cannot be answered. Please restate the question, giving also the total number of positions on the spinner.

What number does each card in a deck of cards represents?

There are different values that are shown on each card. Furthermore, the value of the card depends on the game. In cribbage, for example, Jacks are worth 5. In blackjack, they are worth 10.

What is the function of an antacid?

Antacids neutralize acid, especially acid found in the stomach. They help cause rapid relief for heartburn.

Probability of full house in poker?

The probability of being dealt a full house in poker is 3744/2,598,960 or 1 in 694. The probability of moving up from a full house in poker is 4/47.

Check with http://www.indepthinfo.com/probability-poker/full-house.shtml

The average number of days school is closed due to snow during the winter in a certain city in the eastern part of US is 4.what is the probability that the school in this city will close?

The probability of the school in this city being closed is the same as the probability of Iceland being covered in a cloud of ash to the extent of Airplanes being cancelled.....so yea, pretty good :D

Limitations of statistics?

One limitation of statistics is that everyone doesnÕt understand them or how to use them. There is also a possibility for errors.

Is Statistically independent random variables are uncorrelated and vice versa?

the statistically independent random variables are uncorrelated but the converse is not true ,i want a counter example,

A coin is tossed 5 times find the probability that the 6th toss is a tail given tht the first 5 tosses were all tails?

The chance is 50%-50% that it will be heads or tails; this does not change regardless of the number of previous tosses and their results.

What are the 13 cards in each suit in a deck of 52 playing cards?

The 13 cards in each suit in a deck of 52 playing cards are: Ace, 2, 3, 4, 5, 6, 7, 8, 9, 10, Jack, Queen & King.

Trending Questions
What is the probability that Alex makes both of his shots? What is the probability of spinning a 6 on a 6 number spinner with to number 6? How is experimental probability used to make predictions? What is the probability of getting a even number in a deck of cards? What is the probability of a face card given that the card is black? Taylor tossed one green number cube and one red number cube each cube is numbered from 1-6what is the probility that the sum of the numbers on the cube is 4 write the answer as a fraction in simplest? What is the probability of getting a queen and a heart? How many sides are there in a die? Can the number 0.56 be the probability of an event? A tires life is 100 000km You buy 4 tires What is the probability that all 4 tires will wear out before 100 000km? What is conditional expectation of 2 uncorrelated random variables-is it equal to unconditional expectation? What does mutually exclusive events mean? Is the probability of two independent events occurring together the same as the probability of each occurring alone? Given a Poisson distribution with mean equals 2 Find P x5? What heads is on a 14094766? How many different 5-card hands can be dealt from a standard deck of 52 cards? What is the probilitly of rolling a single die to a even number or a number less than six? Whats the difference between random errors and systematic errors? If you throw a 6 sided die 96 times how many times would it land on a factor of 24? What does z mean on necklace?